Greg spent $36.64 on $0.80 and $0.12 stamps. He bought a total of 107 stamps. How many of the $0.12 stamps did he buy?
Answer by macston(5194)

.
E=number of $0.80 stamps; T=number of $0.12 stamps
.
E+T=107
E=107-T
.
$0.80E+$0.12T=$36.64
$0.80(107-T)+$0.12T=$36.64
$85.60-$0.80T+$0.12T=$36.64
-$0.68T=-$48.96
T=72
.
ANSWER: He bought seventy-two $0.12 stamps.
.
CHECK:
E=107-T=107-72=35
He bought 35 $0.80 stamps.
.
$0.80E+$0.12T=$36.64
$0.80(35)+$0.12(72)=$36.64
$28.00+$8.64=$36.64
$36.64=$36.64
.